# LLM4Drive: A Survey of Large Language Models for Autonomous Driving

### v3

@@ 3.2 Perception
Large language models have demonstrated their unique value and strong capabilities in “perception” tasks Radford et al. (2021); Li et al. (2022b, 2023a, 2023b, a). Especially in environments where data is relatively scarce, these models can rely on their few-shot learning characteristics to achieve fast and accurate learning and reasoning P et al. (2023); Lin et al. (2023). This learning ability is of significance in the perception stage of the autonomous driving system, which greatly improves the system’s adaptability and generalization capabilities in changing and complex driving environments. PromptTrack Wu et al. (2023) fuses cross-modal features in a prompt reasoning branch to predict 3D objects. It uses language prompts as semantic cues and combines LLMs with 3D detection tasks and tracking tasks. Although it achieves better performance compared to other methods, the advantages of LLMs do not directly affect the tracking task. Rather, the tracking task serves as a query to assist LLMs in performing 3D detection tasks. HiLM-D Ding et al. (2023) incorporates high-resolution information into multimodal large language models for the Risk Object Localization and Intention and Suggestion Prediction (ROLISP) task. It combines LLMs with 2D detection tasks and obtains better performance in detection tasks and QA tasks compared to other multi-modal large models such as Video-LLaMa Zhang et al. (2023a), eP-ALM Shukor et al. (2023). It is worth noting to point out one potential limitation of the dataset: each video contains only one risk object, which might not capture the complexity of real-world scenarios. Keysan et al. (2023) integrates pre-trained language models as text-based input encoders for the autonomous driving trajectory prediction task. Joint encoders(image and text) over both modalities perform better than using a single encoder in isolation. While the joint model significantly improves the baseline, its performance has not reached the state-of-the-art level yet Deo et al. (2021); Gilles et al. (2021).
@@
PromptTrack Wu et al. (2023) uses the Average Multiple Object Tracking Precision (AMOTA) metric Bernardin and Stiefelhagen (2008), the Average MultiObject Tracking Precision (AMOTP) Bashar et al. (2022) and Identity Switches (IDS) Huang et al. (2023b) metrics. HiLM-D Ding et al. (2023) uses the BLEU-4 Papineni et al. (2002), METEOR Banerjee and Lavie (2005), CIDERVedantam et al. (2015) and SPICE Anderson et al. (2016), IoU Rezatofighi et al. (2019) as metrics to compare with the state-of-the-art.  Keysan et al. (2023) uses the standard evaluation metrics that are provided in the nuScenes-devkit  Caesar et al. (2019); Fong et al. (2021): minimum Average Displacement Error (minADEk), Final Displacement Error (minFDEk), and the miss rate over 2 meters.
@@ 3.3 Question Answering
Question-Answering is an important task that has a wide range of applications in intelligent transportation, assisted driving, and autonomous vehicles Xu et al. (2021a, b). It mainly reflects through different question and answer paradigms, including traditional QA mechanism Tang et al. (2023) and more detailed visual QA methods Xu et al. (2023). Tang et al. (2023) constructs the domain knowledge ontology by “chatting” with ChatGPT. It develops a web-based assistant to enable manual supervision and early intervention at runtime and it guarantees the quality of fully automated distillation results. This question-and-answer system enhances the interactivity of the vehicle, transforms the traditional one-way human-machine interface into an interactive communication experience, and might be able to cultivate the user’s sense of participation and control. These sophisticated models Tang et al. (2023); Xu et al. (2023), equipped with the ability to parse, understand, and generate human-like responses, are pivotal in real-time information processing and provision. They design comprehensive questions related to the scene, including but not limited to vehicle states, navigation assistance, and understanding of traffic situations. Yang et al. (2023b) provides a human-centered perspective and gives several key insights through different prompt designs to enable LLMs to achieve AD system requirements within the cabin. Dolphins Ma et al. (2023a) enhances reasoning capabilities through the innovative Grounded Chain of Thought (GCoT) process and specifically adapts to the driving domain by building driving-specific command data and command adjustments. LingoQA Marcu et al. (2023) develops a QA benchmark and datasets, details are in 3.5 and 4.
@@
In terms of QA tasks, NLP’s metric is often used. In DriveFPT4 Xu et al. (2023), it uses BLEU-4 Papineni et al. (2002), METEOR Banerjee and Lavie (2005), CIDERVedantam et al. (2015) and chatgpt score Fu et al. (2023b). In Yang et al. (2023b), it adapts accuracy at the individual question level and the command level which includes some sub-questions.
@@ 3.4 Generation
The common generative models include the Variational Auto-Encoder(VAE) Kingma and Welling (2022), Generative Adversarial Network(GAN) Goodfellow et al. (2014), Normalizing Flow(Flow)Rezende and Mohamed (2016), and Denoising Diffusion Probabilistic Model(Diffusion)Ho et al. (2020). With diffusion models have recently achieved great success in text-to-image Ronneberger et al. (2015); Rombach et al. (2021); Ramesh et al. (2022), some research has begun to study using diffusion models to generate autonomous driving images or videos. DriveDreamer Wang et al. (2023d) is a world model derived from real-world driving scenarios. It uses text, initial image, HDmap, and 3Dbox as input, then generates high-quality driving videos and reasonable driving policies. Similarly, Driving Diffusion Li et al. (2023c) adopts a 3D layout as a control signal to generate realistic multi-view videos. GAIA-1 Hu et al. (2023a) leverages video, text, and action inputs to generate traffic scenarios, environmental elements, and potential risks. In these methods, text encoder both adopt CLIP Radford et al. (2021) which has a better alignment between image and text. In addition to generating autonomous driving videos, traffic scenes can also be generated. CTG++ Zhong et al. (2023) is a scene-level diffusion model that can generate realistic and controllable traffic. It leverages LLMs for translating a user query into a differentiable loss function and use a diffusion model to transform the loss function into realistic, query compliant trajectories. MagicDrive Gao et al. (2023) generates highly realistic images, exploiting geometric information from 3D annotations by independently encoding road maps, object boxes, and camera parameters for precise, geometry-guided synthesis. This approach effectively solves the challenge of multi-camera view consistency. Although it achieves better performance in terms of generation fidelity compared to BEVGen Swerdlow et al. (2023) and BEVControl Yang et al. (2023a), it also faces huge challenges in some complex scenes, such as night views and unseen weather conditions. ADriver-I Jia et al. (2023a) combines Multimodal Large Language Models(MLLM) and Video Diffusion Model(VDM) to predict the control signal of current frame and the future frames. It shows impressive performance on nuScenes and their private datasets. However, MLLM and VDM are trained separately, which fails to optimize jointly. Driving into the Future  Wang et al. (2023f) develops a multiview world model, named Drive-WM, which is capable of generating high-quality, controllable, and consistent multi-view videos in autonomous driving scenes. It explores the potential application of the world model in end-to-end planning for autonomous driving.
These methods explore the customized authentic generations of autonomous driving data. Although these diffusion-based models achieved good results on video and image-generated metrics, it is still unclear whether they could really be used in closed-loop to really boost the performance of the autonomous driving system.
@@
DriveDreamer Wang et al. (2023d) and DrivingDiffusion Li et al. (2023c) use the frame-wise Frechet Inception Distance (FID) Parmar et al. (2022) to evaluate the quality of generated images and the Frechet Video Distance (FVD) Unterthiner et al. (2019) for video quality evaluation. DrivingDiffusion also uses average intersection crossing (mIoU) Rezatofighi et al. (2019) scores for drivable areas and NDS Yin et al. (2021) for all the object classes by comparing the predicted layout with the ground-truth BEV layout. CTG++ Zhong et al. (2023) following Xu et al. (2022); Zhong et al. (2022), uses the failure rate, Wasserstein distance between normalized histograms of driving profiles, realism deviation (real), and scene-level realism metric (rel real) as metrics. MagicDrive Gao et al. (2023) utilizes segmentation metrics such as Road mIoU and Vehicle mIoU Taran et al. (2018), as well as 3D object detection metrics like mAPHenderson and Ferrari (2017) and NDS Yin et al. (2021). ADriver-I Jia et al. (2023a) adapts L1 error of speed and steering angle of the current frame, Frechet Inception Distance(FID), and Frechet Video Distance(FVD) as evaluation indicators.
@@ 3.5 Evaluation & Benchmark
In terms of benchmark, LMDrive Shao et al. (2023) introduces LangAuto(Language-guided Autonomous Driving) CARLA benchmark. It covers various driving scenarios in 8 towns and takes into account 16 different environmental conditions. It contains three tracks: LangAuto track (updates navigation instructions based on location and is divided into sub-tracks of different route lengths), LangAuto-Notice track (adds notification instructions based on navigation instructions), and LangAuto-Sequential track (Combining consecutive instructions into a single long instruction). In addition, LangAuto also uses three main evaluation indicators: route completion, violation score, and driving score to comprehensively evaluate the autonomous driving system’s ability to follow instructions and driving safety. LingoQA Marcu et al. (2023) developed LingoQA which is used for evaluating video question-answering models for autonomous driving. The evaluation system consists of three main parts: a GPT-4-based evaluation, which determines whether the model’s answers are consistent with human answers; and the Lingo-Judge metric, which evaluates the model using a trained text classifier called Lingo-Judge Accuracy of answers; and correlation analysis with human ratings. This analysis involves multiple human annotators rating responses to 17 different models on a scale of 0 to 1, which are interpreted as the likelihood that the response accurately solves the problem. Reason2Drive Nie et al. (2023) introduces the protocol to measure the correctness of the reasoning chains to resolve semantic ambiguities. The evaluation process includes four key metrics: Reasoning Alignment, which measures the extent of overlap in logical reasoning; Redundancy, aimed at identifying any repetitive steps; Missing Step, which focuses on pinpointing any crucial steps that are absent but necessary for problem-solving; and Strict Reason, which evaluates scenarios involving visual elements. LaMPilot Ma et al. (2023b) is an innovative benchmark test used to evaluate the instruction execution capabilities of autonomous vehicles, including three parts: a simulator, a data set, and an evaluator. It employs Python Language Model Programs (LMPs) to interpret human-annotated driving instructions and execute and evaluate them within its framework. Tanahashi et al. (2023) evaluated the two core capabilities of large language models (LLMs) in the field of autonomous driving: first, the spatial awareness decision-making ability, that is, LLMs can accurately identify the spatial layout based on coordinate information; second, the ability to follow traffic rules to ensure that LLMs Ability to strictly abide by traffic laws while driving.

### v4

@@ 3.2 Perception
Large language models have demonstrated their unique value and strong capabilities in “perception” tasks Radford et al. (2021); Li et al. (2022b, 2023a, 2023b, a). Especially in environments where data is relatively scarce, these models can rely on their few-shot learning characteristics to achieve fast and accurate learning and reasoning P et al. (2023); Lin et al. (2023). This learning ability is of significance in the perception stage of the autonomous driving system, which greatly improves the system’s adaptability and generalization capabilities in changing and complex driving environments. PromptTrack Wu et al. (2023) fuses cross-modal features in a prompt reasoning branch to predict 3D objects. It uses language prompts as semantic cues and combines LLMs with 3D detection tasks and tracking tasks. Although it achieves better performance compared to other methods, the advantages of LLMs do not directly affect the tracking task. Rather, the tracking task serves as a query to assist LLMs in performing 3D detection tasks. HiLM-D Ding et al. (2023) incorporates high-resolution information into multimodal large language models for the Risk Object Localization and Intention and Suggestion Prediction (ROLISP) task. It combines LLMs with 2D detection tasks and obtains better performance in detection tasks and QA tasks compared to other multi-modal large models such as Video-LLaMa Zhang et al. (2023a), eP-ALM Shukor et al. (2023). It is worth noting to point out one potential limitation of the dataset: each video contains only one risk object, which might not capture the complexity of real-world scenarios. Keysan et al. (2023) integrates pre-trained language models as text-based input encoders for the autonomous driving trajectory prediction task. Joint encoders(image and text) over both modalities perform better than using a single encoder in isolation. While the joint model significantly improves the baseline, its performance has not reached the state-of-the-art level yet Deo et al. (2021); Gilles et al. (2021). LC-LLM Peng et al. (2024) is designed for lane change prediction, leveraging LLM capabilities to understand complex scenarios, enhancing prediction performance, and providing explainable predictions by generating explanations for lane change intentions and trajectories. AIDE Liang et al. (2024) introduces a paradigm for an automatic data engine, incorporating automatic data querying and labeling using VLM, and continual learning with pseudo labels. It introduces a new benchmark to evaluate such automated data engines for AV perception that allows combined insights across multiple paradigms of open vocabulary detection, semi-supervised, and continual learning. Context-aware Motion Prediction Zheng et al. (2024a) designs and conducts prompt engineering to enable GPT4-V to comprehend complex traffic scenarios. It combines the context information outputted by GPT4-V with MTR Shi et al. (2023) to enhance motion prediction.
@@
PromptTrack Wu et al. (2023) uses the Average Multiple Object Tracking Precision (AMOTA) metric Bernardin and Stiefelhagen (2008), the Average MultiObject Tracking Precision (AMOTP) Bashar et al. (2022) and Identity Switches (IDS) Huang et al. (2023b) metrics. HiLM-D Ding et al. (2023) uses the BLEU-4 Papineni et al. (2002), METEOR Banerjee and Lavie (2005), CIDERVedantam et al. (2015) and SPICE Anderson et al. (2016), IoU Rezatofighi et al. (2019) as metrics to compare with the state-of-the-art.  Keysan et al. (2023) uses the standard evaluation metrics that are provided in the nuScenes-devkit  Caesar et al. (2019); Fong et al. (2021): minimum Average Displacement Error (minADEk), Final Displacement Error (minFDEk), and the miss rate over 2 meters. LC-LLM Peng et al. (2024) uses RMSE (lat) to assess lateral and longitudinal prediction error. LeGo-Drive Paul et al. (2024) adopts minFDE (minimum Final Displacement Error) and L2 distance between the goal location and the trajectory endpoint. as evaluation metrics. Context-aware Motion Prediction Zheng et al. (2024a) use the mean Average Precision (mAP) in official WOMD Ettinger et al. (2021) evaluation.
@@
NuInstruct
Ding et al. (2024)
Perception
Prediction
VQA
91k QA pairs, 17 subtasks
GPT-4
Human
Integrates multi-view information,
requiring responses from multiple
perspectives, with balanced view
distribution for perception tasks.
OpenDV-2K
Yang et al. (2024)
Perception
Prediction
VQA
2059 hours of videos paired
with texts(1747 hours from
YouTube and 312 hours from
public datasets).
BLIP-2
A large-scale multimodal dataset
for autonomous driving, to support
the training of a generalized video
prediction model.
@@ 3.3 Question Answering
Question-Answering is an important task that has a wide range of applications in intelligent transportation, assisted driving, and autonomous vehicles Xu et al. (2021a, b). It mainly reflects through different question and answer paradigms, including traditional QA mechanism Tang et al. (2023) and more detailed visual QA methods Xu et al. (2023). Tang et al. (2023) constructs the domain knowledge ontology by “chatting” with ChatGPT. It develops a web-based assistant to enable manual supervision and early intervention at runtime and it guarantees the quality of fully automated distillation results. This question-and-answer system enhances the interactivity of the vehicle, transforms the traditional one-way human-machine interface into an interactive communication experience, and might be able to cultivate the user’s sense of participation and control. These sophisticated models Tang et al. (2023); Xu et al. (2023), equipped with the ability to parse, understand, and generate human-like responses, are pivotal in real-time information processing and provision. They design comprehensive questions related to the scene, including but not limited to vehicle states, navigation assistance, and understanding of traffic situations. Yang et al. (2023b) provides a human-centered perspective and gives several key insights through different prompt designs to enable LLMs to achieve AD system requirements within the cabin. Dolphins Ma et al. (2023a) enhances reasoning capabilities through the innovative Grounded Chain of Thought (GCoT) process and specifically adapts to the driving domain by building driving-specific command data and command adjustments. LingoQA Marcu et al. (2023) develops a QA benchmark and datasets, details are in 3.5 and 4. EM-VLM4AD Gopalkrishnan et al. (2024) is an efficient, lightweight, multi-frame vision language model for Visual Question Answering in autonomous driving, and it only requires much less memory and floating point operations than DriveLM Sima et al. (2023).  Nouri et al. (2024) proposes a prototype of a pipeline of prompts and LLMs that receives an item definition and outputs solutions in the form of safety requirements. Hybrid Reasoning Azarafza et al. (2024) uses Large Language Models (LLMs) with inputs from image-detected objects and sensor data, including parameters like object distance, car speed, direction, and location, to generate precise brake and speed control values based on weather conditions. TransGPT Wang et al. (2024c) is a novel large language model for the transportation domain that comes in two variants—TransGPT-SM for single-modal data and TransGPT-MM for multi-modal data—designed to enhance traffic analysis and modeling by generating synthetic traffic scenarios, explaining traffic phenomena, answering traffic-related questions, offering recommendations, and creating comprehensive traffic reports.
@@
In terms of QA tasks, NLP’s metric is often used. In DriveFPT4 Xu et al. (2023) and EM-VLM4AD Gopalkrishnan et al. (2024), it uses BLEU-4 Papineni et al. (2002), METEOR Banerjee and Lavie (2005), CIDERVedantam et al. (2015) and chatgpt score Fu et al. (2023b). In Yang et al. (2023b), it adapts accuracy at the individual question level and the command level which includes some sub-questions.
@@ 3.4 Generation
The common generative models include the Variational Auto-Encoder(VAE) Kingma and Welling (2022), Generative Adversarial Network(GAN) Goodfellow et al. (2014), Normalizing Flow(Flow)Rezende and Mohamed (2016), and Denoising Diffusion Probabilistic Model(Diffusion)Ho et al. (2020). With diffusion models have recently achieved great success in text-to-image Ronneberger et al. (2015); Rombach et al. (2021); Ramesh et al. (2022), some research has begun to study using diffusion models to generate autonomous driving images or videos. DriveDreamer Wang et al. (2023e) is a world model derived from real-world driving scenarios. It uses text, initial image, HDmap, and 3Dbox as input, then generates high-quality driving videos and reasonable driving policies. Similarly, Driving Diffusion Li et al. (2023c) adopts a 3D layout as a control signal to generate realistic multi-view videos. GAIA-1 Hu et al. (2023a) leverages video, text, and action inputs to generate traffic scenarios, environmental elements, and potential risks. In these methods, text encoder both adopt CLIP Radford et al. (2021) which has a better alignment between image and text. In addition to generating autonomous driving videos, traffic scenes can also be generated. CTG++ Zhong et al. (2023) is a scene-level diffusion model that can generate realistic and controllable traffic. It leverages LLMs for translating a user query into a differentiable loss function and use a diffusion model to transform the loss function into realistic, query compliant trajectories. MagicDrive Gao et al. (2023) generates highly realistic images, exploiting geometric information from 3D annotations by independently encoding road maps, object boxes, and camera parameters for precise, geometry-guided synthesis. This approach effectively solves the challenge of multi-camera view consistency. Although it achieves better performance in terms of generation fidelity compared to BEVGen Swerdlow et al. (2023) and BEVControl Yang et al. (2023a), it also faces huge challenges in some complex scenes, such as night views and unseen weather conditions. ADriver-I Jia et al. (2023a) combines Multimodal Large Language Models(MLLM) and Video Diffusion Model(VDM) to predict the control signal of current frame and the future frames. It shows impressive performance on nuScenes and their private datasets. However, MLLM and VDM are trained separately, which fails to optimize jointly. Driving into the Future  Wang et al. (2023g) develops a multiview world model, named Drive-WM, which is capable of generating high-quality, controllable, and consistent multi-view videos in autonomous driving scenes. It explores the potential application of the world model in end-to-end planning for autonomous driving. ChatScene Zhang et al. (2024) designs an LLM-based agent that generates and simulates challenging safety-critical scenarios in CARLA, improving the collision avoidance capabilities and robustness of autonomous vehicles. REvolve Hazra et al. (2024) is an evolutionary framework utilizing GPT-4 to generate and refine reward functions for autonomous driving through human feedback. The reward function is used for RL, and the score is achieved closely by human driving standards. GenAD Yang et al. (2024) is a large-scale video prediction model for autonomous driving that uses extensive web-sourced data and novel temporal reasoning blocks to handle diverse driving scenarios, generalize to unseen datasets in a zero-shot manner, and adapt for action-conditioned prediction or motion planning. DriveDreamer-2 Zhao et al. (2024) builds on DriveDreamer with a Large Language Model (LLM), generates customized and high-quality multi-view driving videos by converting user queries into agent trajectories and HDMaps, enhancing training for driving perception methods. ChatSim Wei et al. (2024) enable editable photo-realistic 3D driving scene simulations via natural language commands with external digital assets, leverages a large language model agent collaboration framework and novel multi-camera neural radiance field and lighting estimation methods to produce scene-consistent, high-quality outputs. LLM-Assisted Light Wang et al. (2024a) integrates the human-mimetic reasoning capabilities of LLMs, enabling the signal control algorithm to interpret and respond to complex traffic scenarios with the nuanced judgment typical of human cognition. It developed a closed-loop traffic signal control system, integrating LLMs with a comprehensive suite of interoperable tools. LangProp Ishida et al. (2024) is a framework that iteratively optimizes code generated by large language models (LLMs) using both supervised and reinforcement learning, automatically evaluating code performance, catching exceptions, and feeding results back to the LLM to improve code generation for autonomous driving in CARLA. These methods explore the customized authentic generations of autonomous driving data. Although these diffusion-based models achieved good results on video and image-generated metrics, it is still unclear whether they could really be used in closed-loop to really boost the performance of the autonomous driving system.
@@
DriveDreamer Wang et al. (2023e), DriveDreamer-2 Zhao et al. (2024), DrivingDiffusion Li et al. (2023c) and GenAD Yang et al. (2024) use the frame-wise Frechet Inception Distance (FID) Parmar et al. (2022) to evaluate the quality of generated images and the Frechet Video Distance (FVD) Unterthiner et al. (2019) for video quality evaluation. DrivingDiffusion also uses average intersection crossing (mIoU) Rezatofighi et al. (2019) scores for drivable areas and NDS Yin et al. (2021) for all the object classes by comparing the predicted layout with the ground-truth BEV layout. CTG++ Zhong et al. (2023) following Xu et al. (2022); Zhong et al. (2022), uses the failure rate, Wasserstein distance between normalized histograms of driving profiles, realism deviation (real), and scene-level realism metric (rel real) as metrics. MagicDrive Gao et al. (2023) utilizes segmentation metrics such as Road mIoU and Vehicle mIoU Taran et al. (2018), as well as 3D object detection metrics like mAPHenderson and Ferrari (2017) and NDS Yin et al. (2021). ADriver-I Jia et al. (2023a) adapts L1 error of speed and steering angle of the current frame, Frechet Inception Distance(FID), and Frechet Video Distance(FVD) as evaluation indicators. ChatScene Zhang et al. (2024) provide a thorough evaluation of various scenario generation algorithms. These are assessed based on the collision rate (CR), overall score (OS), and average displacement error (ADE). REvolve Hazra et al. (2024) adopts fitness scores and episodic steps as metrics in AirSim simulator Shah et al. (2017).
@@ 3.5 Evaluation & Benchmark
In terms of benchmark, LMDrive Shao et al. (2023) introduces LangAuto(Language-guided Autonomous Driving) CARLA benchmark. It covers various driving scenarios in 8 towns and takes into account 16 different environmental conditions. It contains three tracks: LangAuto track (updates navigation instructions based on location and is divided into sub-tracks of different route lengths), LangAuto-Notice track (adds notification instructions based on navigation instructions), and LangAuto-Sequential track (Combining consecutive instructions into a single long instruction). In addition, LangAuto also uses three main evaluation indicators: route completion, violation score, and driving score to comprehensively evaluate the autonomous driving system’s ability to follow instructions and driving safety. LingoQA Marcu et al. (2023) developed LingoQA which is used for evaluating video question-answering models for autonomous driving. The evaluation system consists of three main parts: a GPT-4-based evaluation, which determines whether the model’s answers are consistent with human answers; and the Lingo-Judge metric, which evaluates the model using a trained text classifier called Lingo-Judge Accuracy of answers; and correlation analysis with human ratings. This analysis involves multiple human annotators rating responses to 17 different models on a scale of 0 to 1, which are interpreted as the likelihood that the response accurately solves the problem. Reason2Drive Nie et al. (2023) introduces the protocol to measure the correctness of the reasoning chains to resolve semantic ambiguities. The evaluation process includes four key metrics: Reasoning Alignment, which measures the extent of overlap in logical reasoning; Redundancy, aimed at identifying any repetitive steps; Missing Step, which focuses on pinpointing any crucial steps that are absent but necessary for problem-solving; and Strict Reason, which evaluates scenarios involving visual elements. LaMPilot Ma et al. (2023b) is an benchmark test used to evaluate the instruction execution capabilities of autonomous vehicles, including three parts: a simulator, a data set, and an evaluator. It employs Python Language Model Programs (LMPs) to interpret human-annotated driving instructions and execute and evaluate them within its framework. Tanahashi et al. (2023) evaluated the two core capabilities of large language models (LLMs) in the field of autonomous driving: first, the spatial awareness decision-making ability, that is, LLMs can accurately identify the spatial layout based on coordinate information; second, the ability to follow traffic rules to ensure that LLMs Ability to strictly abide by traffic laws while driving.  Tang et al. (2024) tests three OpenAI LLMs and several other LLMs on UK Driving Theory Test Practice Questions and Answers, and only GPT-4o passed the test, indicating that the performance of LLMs still needs to be further improved.  Kong et al. (2024) has developed an LLM-based safe autonomous driving framework, which evaluates and enhances the performance of existing LLM-AD methods in driving safety, sensitive data usage, token consumption and alignment scenarios by integrating security assessment agents. DriveSim Sreeram et al. (2024) is a specialized simulator that creates diverse driving scenarios to test and benchmark MLLMs’ understanding and reasoning of real-world driving scenes from a fixed in-car camera perspective. OmniDrive Wang et al. (2024d) introduces a comprehensive benchmark for visual question-answering tasks related to 3D driving, ensuring strong alignment between agent models and driving tasks through scene description, traffic regulation, 3D grounding, counterfactual reasoning, decision-making, and planning. AIDE Liang et al. (2024) proposes an automatic data engine design paradigm, which features automatic data query and labeling using VLM and continuous learning with pseudo-labels. It also introduces a new benchmark to evaluate such automatic data engines for self-driving car perception, providing comprehensive insights across multiple paradigms including open vocabulary detection, semi-supervised learning, and continuous learning. ELM Zhou et al. (2024) is proposed to understand driving scenes over long-scope space and time, showing promising generalization performance in handling complex driving scenarios. LimSim++ Fu et al. (2024) introduces an open-source evaluation platform for (M)LLM in autonomous driving, supporting scenario understanding, decision-making, and evaluation systems.